# Vim-F: Visual State Space Model Benefiting from Learning in the Frequency Domain

### v1

@@ Abstract
In recent years, State Space Models (SSMs) with efficient hardware-aware designs, known as the Mamba deep learning models, have made significant progress in modeling long sequences such as language understanding. Therefore, building efficient and general-purpose visual backbones based on SSMs is a promising direction. Compared to traditional convolutional neural networks (CNNs) and Vision Transformers (ViTs), the performance of Vision Mamba (ViM) methods is not yet fully competitive. To enable SSMs to process image data, ViMs typically flatten 2D images into 1D sequences, inevitably ignoring some 2D local dependencies, thereby weakening the model’s ability to interpret spatial relationships from a global perspective. We use Fast Fourier Transform (FFT) to obtain the spectrum of the feature map and add it to the original feature map, enabling ViM to model a unified visual representation in both frequency and spatial domains. The introduction of frequency domain information enables ViM to have a global receptive field during scanning. We propose a novel model called Vim-F, which employs pure Mamba encoders and scans in both the frequency and spatial domains. Moreover, we question the necessity of position embedding in ViM and remove it accordingly in Vim-F, which helps to fully utilize the efficient long-sequence modeling capability of ViM. Finally, we redesign a patch embedding for Vim-F, leveraging a convolutional stem to capture more local correlations, further improving the performance of Vim-F. Code is available at: https://github.com/yws-wxs/Vim-F.
1 Introduction
@@
Visual representation learning stands as a pivotal research topic within the realm of computer vision. In the early era of deep learning, Convolutional Neural Networks (CNNs) occupied a dominant position. CNNs have several built-in inductive biases that make them well-suited to a wide variety of computer vision applications. In addition, CNNs are inherently efficient because when used in a sliding-window manner, the computations are shared. Inspired by the significant success of the Transformer architecture in the field of Natural Language Processing (NLP), researchers have recently applied Transformer to computer vision tasks. Except for the initial "patch embedding" module, which splits an image into a sequence of patches, Vision Transformers (ViTs) do not introduce image-specific inductive biases and follow the original NLP Transformers design concept as much as possible. Compared to CNNs, ViTs typically exhibit superior performance, which can be attributed to the global receptive field and dynamic weights facilitated by the attention mechanism. However, the attention mechanism requires quadratic complexity in terms of image sizes, resulting in significant computational cost when addressing downstream dense prediction tasks, such as object detection and instance segmentation.
@@
Derived from the classic Kalman filter model, modern State Space Models (SSMs) excel at capturing long-range dependencies and reap the benefits of parallel training. The Visual Mamba Models (ViMs) Zhu et al., (2024); Liu et al., (2024); Huang et al., (2024); Pei et al., (2024), which are inspired by recently proposed SSMs Gu and Dao, (2023); Mehta et al., (2023), utilize the Selective Scan Space State Sequential Model (S6) to compress previously scanned information into hidden states, effectively reducing quadratic complexity to linear. Most of these studies integrate the original SSM structure of Mamba into their basic models, thereby necessitating the conversion of 2D images into 1D sequences for SSM-based processing. However, due to the non-causal nature of visual data, flattening spatial data into one-dimensional tokens destroys local two-dimensional dependencies, impairing the model’s capacity to accurately interpret spatial relationships. Vim Zhu et al., (2024) addresses this issue by scanning in bidirectional horizontal directions, while VMamba Liu et al., (2024) adds vertical scanning, enabling each element in the feature map to integrate information from other locations in different directions. Subsequent works, such as LocalMamba Huang et al., (2024) and EfficientVMamba Pei et al., (2024), have designed a series of novel scanning strategies in the spatial domain, aimed at capturing local dependencies while maintaining a global view. We believe that these strategies do not significantly enhance the model’s ability to obtain a better global receptive field. These efforts have prompted us to focus on the value of scanning in the frequency domain. We know that the use of 2D Discrete Fourier transform (DFT) to convert feature maps into spectrograms does not alter their shapes. The value $P(x,y)$ at any point $(x,y)$ in the spectrogram depends on the entire original feature map. Therefore, scanning in the frequency domain ensures that the model always has a good global receptive field. On the other hand, without considering spectral shifts, the low-frequency components are located in the corners of the spectrogram, whereas the high-frequency components are centered. Scanning across the spectrogram typically alternates between accessing low and high frequencies, which may facilitate balanced modeling. Therefore, we adopt the most direct solution, which is to directly add the feature map with its amplitude spectrum, and then let the Vision Mamba encoder learn the fused semantic information. Using the Fast Fourier Transform (FFT) algorithm to calculate the amplitude spectrum not only introduces no trainable parameters but also has minimal impact on performance. To clearly demonstrate the effectiveness and ease of use of our work, we chose Vim as the base model, which to our knowledge is the first pure-SSM-based visual model. Based on the open-source code provided by the authors, our method, called Vim-F, can be easily implemented by adding three lines of code. Based on Vim-F, we improve the patch embedding to capture more local dependencies. Many successful lightweight transformer models, such as EfficientFormer Li et al., (2022) and MobileViT Mehta and Rastegari, (2022), start their networks with a convolutional stem as patch embedding. This is attributed to the inherent spatial inductive bias of CNN, which is beneficial for capturing local information. In addition, CNNs are less sensitive to data augmentation, and the training process of a hybrid architecture is more stable. Inspired by these works, we design a patch embedding for Vim-F. Specifically, we use multiple convolutional layers for downsampling and then flatten the feature maps. Vim-F that employs this patch embedding module is referred to as Vim-F(H). Finally, we believe that the Mamba model is essentially still a recurrent neural network (RNN), so Vim lacks justification for using position embedding, but removing it will significantly reduce Vim’s performance. Conversely, the use of position embedding has less impact on the performance of Vim-F(H), so our Vim-F(H) does not use position embedding. We outline our contributions below:
@@
  • Based on Vim Zhu et al., (2024), we propose Vim-F, which enables SSM to scan both in the spatial and frequency domains, ensuring that SSM always has a global view during the scanning process, thereby enhancing the model’s ability to interpret spatial relationships. To the best of our knowledge, our work is the first to propose scanning in the frequency domain.
@@
  • We design a novel patch embedding module that utilizes a mixture of overlapping and non-overlapping convolutions with small kernel sizes to downsample the feature map and then flatten it. The Vim-F with this patch embedding is referred to as Vim-F(H). Additionally, we question the necessity of position embedding in Vim, thus Vim-F(H) does not utilize position embedding.
  • Through extensive experiments on image classification, object detection, and instance segmentation tasks, the simple Vim variants, Vim-F(H), achieve better performance while maintaining comparable computational costs and parameters. Our work demonstrates that the pure Mamba encoder can benefit from scanning in the frequency domain and introducing more 2D local correlation during the image flattening stage.
2 Related works
2.1 Generic vision backbone design
Nowdays, the realm of vision tasks has been predominantly governed by CNNs and ViTs. Notable CNNs such as VGGNet Simonyan and Zisserman, (2014), Inception Szegedy et al., (2015), ResNet He et al., (2016),ResNeXt Xie et al., (2017), DenseNet Huang et al., (2017), MobileNet Howard et al., (2017), EfficientNet Tan and Le, (2019), and RegNet Radosavovic et al., (2020) have focused on diverse aspects of accuracy, efficiency, and scalability, while promoting numerous influential design principles that have gained widespread acceptance. Around the same time, the network design philosophy for natural language processing took a completely different path, with Transformers replacing Recurrent Neural Networks (RNNs) to become the dominant backbone architecture. Despite the significant differences in the tasks of interest between the language and vision domains, it is noteworthy that the two streams unexpectedly converged. This convergence occurred due to the introduction of ViTs Dosovitskiy et al., (2021), which revolutionized the landscape of network architecture design. The biggest challenge for ViT comes from the design of global attention, which has quadratic complexity relative to the input size. Hierarchical Transformers employ a hybrid approach to solve this problem. For example, the “sliding window” strategy, which involves attention within local windows, has been reintroduced to Transformers Liu et al., (2021), allowing them to behave more similarly to CNNs. Moreover, other studies propose to harness the advantages of CNNs, such as introducing convolution operations, designing hybrid architectures by combining CNN and ViT modules Dai et al., (2021); Srinivas et al., (2021). State Space Models (SSMs) exhibit linear or near-linear scaling complexity with sequence length, making them particularly suitable for handling long sequences. Therefore, SSMs have been widely applied to a range of vision tasks. These applications extend from developing generic foundation models to advancing fields in image segmentation Wang et al., (2024); Ruan and Xiang, (2024), demonstrating the model’s adaptability and potential in visual domain.
2.2 Fourier transforms in neural networks
Because ordinary multiplication in the frequency domain corresponds to a convolution in the time domain, Fast Fourier Transforms (FFTs) have previously been used to approximate or speed up computations in Convolutional Neural Networks Pratt et al., (2017), Recurrent Neural Networks Zhang et al., (2018), and Transformers Tamkin et al., (2020). FNetLee-Thorp et al., (2022) replaces the self-attention sublayer in the Transformer encoder with a standard, non-parametric Fourier transform, achieving an excellent trade-off between speed, memory usage, and accuracy. This suggests that the Fourier transform can serve as an alternative mixing mechanism for hidden representations. GFNet Rao et al., (2021) models the interactions between tokens as a set of learnable global filters that are applied to the spectrum of input features to learn long-range spatial dependencies in the frequency domain with log-linear complexity.
3 Preliminaries
3.1 State Space Models
@@
3.2 Discretization
The primary objective of discretization is to convert ODEs into discrete functions, allowing the model to align with the sampling frequency of the input signal for more efficient computation. Following the work Gupta et al., (2022), the continuous parameters ($\bm{A}$, $\bm{B}$) can be discretized by zero-order hold rule with a given sample timescale $\Delta\in\mathbb{R}^{N}$:
$\bar{\bm{A}}=e^{\Delta\bm{A}},$ (2)
$\bar{\bm{B}}=(\Delta\bm{A})^{-1}(e^{\Delta\bm{A}}-\bm{I})\cdot\Delta\bm{B},$
$\bar{\bm{C}}=\bm{C},$
$h(t)=\bar{\bm{A}}h(t-1)+\bar{\bm{B}}x(t),$
$y(t)=\bar{\bm{C}}h(t),$
where $\bm{\bar{A}}\in\mathbb{R}^{N\times N}$, $\bm{\bar{B}}\in\mathbb{R}^{D\times N}$ and $\bm{\bar{C}}\in\mathbb{R}^{N\times D}$.
To improve computational efficiency, the iterative process described in equation 2 can be accelerated through parallel computing using global convolutional operations:
$\begin{split}\bm{y}&=\bm{x}\circledast\bm{\overline{K}}\\ \text{with}\quad\bm{\overline{K}}&=(\bm{C}\bm{\overline{B}},\bm{C}\overline{\bm{A}\bm{B}},...,\bm{C}\bm{\overline{A}}^{L-1}\bm{\overline{B}}),\end{split}$ (3)
[…]
3.3 Selective State Space Models (S6)
Mamba Gu and Dao, (2023) improves the performance of SSM by introducing Selective State Space Models (S6), allowing the continuous parameters to vary with the input enhances selective information processing across sequences, which extend the discretization process by selection mechanism:
$\bar{\bm{B}}$ $=s_{\bm{B}}(x),$ (4)
$\bar{\bm{C}}$ $=s_{\bm{C}}(x),$
$\Delta$ $=softplus\big{(}\text{Parameter}+s_{\bm{A}}(x)\big{)},$
where $s_{\bm{B}}(x)$ and $s_{\bm{C}}(x)$ are linear functions that project input $x$ into an N-dimensional space, while $s_{\bm{A}}(x)$ broadens a $D$-dimensional linear projection to the necessary dimensions. Parameter represents a trainable parameter matrix. $softplus$ is a mathematical operation defined as: $softplus(x)=log(1+e^{x})$.
Figure 1: The overview of the Vim model.
3.4 Overview of Vim
As shown in Figure 1, Vim first reshapes the 2D image $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ into a sequence of flattened 2D patches $\mathbf{X}_{p}\in\mathbb{R}^{n\times(p^{2}\cdot c)}$, where $c$ is the number of channels, $(h,w)$ is the resolution of the original image, and $(p,p)$ is the resolution of each image patch. The effective sequence length for Vim is therefore $n=hw/p^{2}$. Then, Vim linearly projects $\mathbf{X}_{p}$ into vectors of size $d$ and adds them to the position embeddings $E_{pos}\in\mathbb{R}^{(n+1)\times d}$, as follows:
@@ 3.4 Overview of Vim
where $\mathbf{Vim}$ is the proposed vision mamba block, $L$ is the number of layers, and $\mathbf{Norm}$ is the normalization layer.
4 Methodology
In this section, we commence with a thorough introduction to the scanning strategy of Vim-F. Subsequently, we present the specific details of the patch embedding in Vim-F(H). Finally, we analyze the efficiency of Vim-F(H). Through these steps, we aim to provide a comprehensive understanding of Vim-F(H).
4.1 2D scanning in the frequency domain
@@ 4 Methodology
As shown in Figure 1, the flattening method used in Vim disrupts these local dependencies and significantly increases the distance between vertically adjacent tokens. We address this limitation by introducing frequency-domain scanning. Specifically, we first perform a Fourier transform on the feature map to obtain:
@@
where $j$ represents the imaginary unit, $H$ and $W$ represent the height and width of the feature map, respectively. $f(x,y)\in\mathbb{R}^{H\times W}$ represents the corresponding value at the coordinate $(x,y)$ in the feature map, and $F(u,v)$ represents the corresponding value at the coordinate $(u,v)$ in the frequency domain. $|F(u,v)|\in\mathbb{R}^{H\times W}$ represents the modulus of $F(u,v)$, which is also known as the amplitude spectrum. Fourier transform has translation invariance. In other words, if a function is translated in the spatial domain, its Fourier transform remains the same as the original function. This property is quite meaningful for addressing the issue of increasing distances between vertically adjacent tokens, as shown in Figure 2. The amplitude spectra of figures (a) and (b) are identical, so scanning in the frequency domain helps reduce the inductive bias introduced by the scanning strategy. Therefore, frequency domain scanning does not require fancy strategies, and it can be consistent with Vim’s original spatial domain scanning strategy. Finally, we use two trainable parameters as the superposition coefficients of the feature map and the corresponding spectrogram to dynamically adjust the information intensity in the spatial and frequency domains. The Mamba encoder establishes a unified representation of visual information in the spatial and frequency domains.
Figure 3: The overall architecture of patch embedding in Vim-F(H).
4.2 Patch embedding
The overall architecture of the patch embedding module is shown in Figure 3.
4.2.1 Convolutional stem
Table 1: Convolutional stem for patch embedding.
@@ 4.2.1 Convolutional stem
Specifically, we first utilize a $7\times 7$ convolutional layer with a stride of 4 to model the local correlation of the input image. Then, we downsample the feature map twice through two downsampling blocks. Each downsampling block consists of a $2\times 2$ convolutional layer with a stride of 2 and a $1\times 1$ convolutional layer. The details of the adaptation with the two base models Zhu et al., (2024) (Vim-Ti and Vim-S) are shown in Table 1 and Figure 3.
4.3 Efficiency analysis
Vim-F requires a 2D discrete Fourier transform (DFT) to be performed on the feature map before scanning, thus introducing additional computation. The Cooley-Tukey FFT algorithm decomposes a DFT of a sequence with length N into two DFTs of sub-sequences with length N/2, significantly reducing the computational complexity. During the execution of the algorithm, the data is divided into two parts based on whether they are odd or even, and the DFTs of each part are computed separately before merging the results. For a 2D-DFT with dimensions $(H,W)$, its complexity is approximately $O(H\cdot W\cdot log(H)\cdot log(W))$. Compared to the overall computational complexity of the model, the computational cost of this part can be negligible.
The patch embedding in Vim can be implemented through a $16\times 16$ convolutional layer with a stride of 16. The convolutional stem for patch embedding in Vim-F(H) comprises more convolutional layers, but the computational cost increases moderately due to the use of smaller convolutional kernels. Considering Vim-Ti as the baseline model, the computational load measured in GFLOPs only increases by 0.035. In summary, there is no significant difference in computational cost between Vim-F(H) and Vim. LocalVimHuang et al., (2024) is one of the representative works that improve the scanning strategy for Vim. In terms of efficiency, LocalVim’s selective scanning branches are twice as many as our Vim-F(H), and it uses a complex spatial and channel attention module (SCAttn) to weight the channels and tokens in the features of each branch. Therefore, our Vim-F(H) achieves a better balance between performance and efficiency.
Table 2: Comparison of different backbones on ImageNet-1K classification. :$\dagger$:† : We are unable to reproduce the results through the source code provided by the author, and one possible reason is the significant difference in hardware and the use of linear scaling rule. The results in () are from the original paper.

### v2

@@ Abstract
In recent years, State Space Models (SSMs) with efficient hardware-aware designs, known as the Mamba deep learning models, have made significant progress in modeling long sequences. Compared to traditional convolutional neural networks (CNNs) and Vision Transformers (ViTs), the performance of Vision Mamba (ViM) methods is not yet fully competitive. To enable SSMs to process image data, ViMs typically flatten 2D images into 1D sequences, inevitably ignoring some 2D local dependencies, thereby weakening the model’s ability to interpret spatial relationships from a global perspective. We believe that the introduction of frequency domain information can enable ViM to achieve a better global receptive field during the scanning process. We propose a novel model called Vim-F, which employs pure Mamba encoders and scans in both the frequency and spatial domains. Moreover, we question the necessity of position embedding in ViM and remove it accordingly in Vim-F, which helps to fully utilize the efficient long-sequence modeling capability of ViM. Finally, we redesign a patch embedding for Vim-F, leveraging a convolutional stem to capture more local correlations, further improving the performance of Vim-F. Code is available at: https://github.com/yws-wxs/Vim-F.
Index Terms:
Vision Mamba, State Space Model, Fast Fourier Transform, Scanning strategy
I Introduction
@@
Visual representation learning stands as a pivotal research topic within the realm of computer vision. In the early era of deep learning, Convolutional Neural Networks (CNNs) occupied a dominant position. CNNs have several built-in inductive biases that make them well-suited to a wide variety of computer vision applications. In addition, CNNs are inherently efficient because when used in a sliding-window manner, the computations are shared. Inspired by the significant success of the Transformer architecture in the field of Natural Language Processing (NLP), researchers have recently applied Transformer to computer vision tasks. Except for the initial ”patch embedding” module, which splits an image into a sequence of patches, Vision Transformers (ViTs) do not introduce image-specific inductive biases and follow the original NLP Transformers design concept as much as possible. Compared to CNNs, ViTs typically exhibit superior performance, which can be attributed to the global receptive field and dynamic weights facilitated by the attention mechanism. However, the attention mechanism requires quadratic complexity in terms of image sizes, resulting in significant computational cost when addressing downstream dense prediction tasks, such as object detection and instance segmentation.
Figure 1: The overview of the Vim model.
@@
Derived from the classic Kalman filter model, modern State Space Models (SSMs) excel at capturing long-range dependencies and reap the benefits of parallel training. The Visual Mamba Models (ViMs) [1, 2, 3, 4], which are inspired by recently proposed SSMs [5, 6], utilize the Selective Scan Space State Sequential Model (S6) to compress previously scanned information into hidden states, effectively reducing quadratic complexity to linear. Most of these studies integrate the original SSM structure of Mamba into their basic models, thereby necessitating the conversion of 2D images into 1D sequences for SSM-based processing. However, due to the non-causal nature of visual data, flattening spatial data into one-dimensional tokens destroys local two-dimensional dependencies, impairing the model’s capacity to accurately interpret spatial relationships. Vim [1] addresses this issue by scanning in bidirectional horizontal directions, while VMamba [2] adds vertical scanning, enabling each element in the feature map to integrate information from other locations in different directions. Subsequent works, such as LocalMamba [3] and EfficientVMamba [4], have designed a series of novel scanning strategies in the spatial domain, aimed at capturing local dependencies while maintaining a global view. We believe that these strategies do not significantly enhance the model’s ability to obtain a better global receptive field. These efforts have prompted us to focus on the value of scanning in the frequency domain. We know that the use of 2D Discrete Fourier transform (DFT) to convert feature maps into spectrograms does not alter their shapes. The value $P(x,y)$ at any point $(x,y)$ in the spectrogram depends on the entire original feature map. Therefore, scanning in the frequency domain ensures that the model always has a good global receptive field. On the other hand, without considering spectral shifts, the high-frequency components are located in the corners of the spectrogram, whereas the low-frequency components are centered. Scanning across the spectrogram typically alternates between accessing low and high frequencies, which may facilitate balanced modeling. Therefore, we adopt the most direct solution, which is to add the amplitude spectrum of the feature map to the original feature map, allowing the visual Mamba encoder to learn the fused semantic information. Using the Fast Fourier Transform (FFT) algorithm to calculate the amplitude spectrum not only introduces no trainable parameters but also has minimal impact on performance. To clearly demonstrate the effectiveness and ease of use of our work, we chose Vim as the base model, which to our knowledge is the first pure-SSM-based visual model. Based on the open-source code provided by the authors, our method, called Vim-F, can be easily implemented by adding a few lines of code. Secondly, we improve the patch embedding to capture more local dependencies. Many successful lightweight transformer models, such as EfficientFormer [7] and MobileViT [8], start their networks with a convolutional stem as patch embedding. This is attributed to the inherent spatial inductive bias of CNN, which is beneficial for capturing local information. In addition, CNNs are less sensitive to data augmentation, and the training process of a hybrid architecture is more stable. Inspired by these works, we design a patch embedding for Vim-F. Specifically, we use multiple convolutional layers for downsampling and then flatten the feature maps. Finally, we argue that the Mamba model is essentially still a recurrent neural network (RNN). Therefore, Vim lacks the rationale for using positional embeddings, and thus our Vim-F does not employ positional embeddings. Additionally, we reduce the model’s sensitivity to the order of input sequences by alternately stacking linear attention blocks and Vim-F blocks, while maintaining overall linear complexity. We refer to this hybrid architecture as Vim-F(H). We outline our contributions below:
@@
  • Based on Vim [1], we propose Vim-F, which enables SSM to scan both in the spatial and frequency domains, ensuring that SSM always has a global view during the scanning process, thereby enhancing the model’s ability to interpret spatial relationships. To the best of our knowledge, our work is the first to propose scanning in the frequency domain.
@@
  • We design a novel patch embedding module that utilizes a mixture of overlapping and non-overlapping convolutions with small kernel sizes to downsample the feature map and then flatten it. Additionally, we question the necessity of position embedding in Vim, thus Vim-F does not utilize position embedding.
  • The experimental results demonstrate that the simple Vim variant, Vim-F, achieves better performance while maintaining similar computational costs and parameters. Additionally, introducing linear attention to reduce Vim-F’s sensitivity to the order of input sequences can further enhance the model’s performance.
II Related Works
II-A State Space Models
@@
II-B Overview of Vim
As mentioned earlier, the effectiveness of our work is demonstrated through the improvement of Vim. Therefore, we provide a brief introduction to the overall architecture of Vim. As shown in Fig. 1, Vim first reshapes the 2D image $\mathbf{X}\in\mathbb{R}^{h\times w\times c}$ into a sequence of flattened 2D patches $\mathbf{X}_{p}\in\mathbb{R}^{n\times(p^{2}\cdot c)}$, where $c$ is the number of channels, $(h,w)$ is the resolution of the original image, and $(p,p)$ is the resolution of each image patch. The effective sequence length for Vim is therefore $n=hw/p^{2}$. Then, Vim linearly projects $\mathbf{X}_{p}$ into vectors of size $d$ and adds them to the position embeddings $E_{pos}\in\mathbb{R}^{(n+1)\times d}$, as follows:
@@ II-B Overview of Vim
where $\mathbf{Vim}$ is the proposed vision mamba block, $L$ is the number of blocks, and $\mathbf{Norm}$ is the normalization layer.
@@
III Rethinking the Scanning Strategy of ViMs
The non-causal nature of two-dimensional spatial patterns in images is inherently inconsistent with the causal processing framework of SSM. Therefore, it is meaningful to explore methods for serializing images. However, the majority of existing work on scanning strategies is typically based on a rather idealized assumption that introducing more or more complex inductive biases regarding the order of one-dimensional sequences can potentially enhance Mamba’s ability to understand local image correlations. Theoretically, regardless of the order, reducing the distance between adjacent tokens in one direction in two-dimensional space inevitably increases the distance between adjacent tokens in another direction, thus these strategies are essentially equivalent. Additionally, in the absence of global information serving as a ”frame of reference,” these sorting methods are essentially disconnected, making it inherently difficult for Mamba to gain additional spatial relationship information by scanning multiple sequences. In practice, works involving scanning strategy research, including Vim [1], VMamba [2], LocalMamba [3], EfficientVMamba [4], and PlainMamba [9], lack clear comparative experiments to verify the necessity and effectiveness of their scanning strategies. Zhu et al. [10] conduct a comprehensive experimental investigation on the impact of mainstream scanning directions and their combinations on semantic segmentation of remotely sensed images. Through extensive experiments on the LoveDA, ISPRS Potsdam, and ISPRS Vaihingen datasets, they demonstrate that no single scanning strategy outperforms others, regardless of their complexity or the number of scanning directions involved. In summary, an effective scanning strategy needs to maintain a global receptive field throughout the scanning process and demonstrate its effectiveness through analysis and comparative experiments.
IV Methodology
In this section, we first elaborate on the significance and benefits of frequency-domain scanning. Subsequently, we present the specific details of patch embedding in Vim-F. Finally, we discuss the composition of the Vim-F and the Vim-F(H) block.
IV-A 2D Scanning in the Frequency Domain
The flattening method used in Vim [1] disrupts these local dependencies and significantly increases the distance between vertically adjacent tokens. We address this limitation by introducing frequency-domain scanning. Specifically, we first perform a Fourier transform on the feature map to obtain:
@@ IV Methodology
where $j$ represents the imaginary unit, $H$ and $W$ represent the height and width of the feature map, respectively. $f(x,y)\in\mathbb{R}^{H\times W}$ represents the corresponding value at the coordinate $(x,y)$ in the feature map, and $F(u,v)$ represents the corresponding value at the coordinate $(u,v)$ in the frequency domain. $|F(u,v)|\in\mathbb{R}^{H\times W}$ represents the modulus of $F(u,v)$, which is also known as the amplitude spectrum. Fourier transform has translation invariance. In other words, if a function is translated in the spatial domain, its Fourier transform remains the same as the original function. This property is quite meaningful for addressing the issue of increasing distances between vertically adjacent tokens, as shown in Fig. 2. The amplitude spectra of figures (a) and (b) are identical, so scanning in the frequency domain helps reduce the inductive bias introduced by the scanning strategy. Therefore, frequency domain scanning does not require fancy strategies, and it can be consistent with Vim’s original spatial domain scanning strategy.
IV-B Patch Embedding
Specifically, we first utilize a $7\times 7$ convolutional layer with a stride of 4 to model the local correlation of the input image. Then, we downsample the feature map twice through two downsampling blocks. Each downsampling block consists of a $2\times 2$ convolutional layer with a stride of 2 and a $1\times 1$ convolutional layer. The details of the adaptation with the two base models [1] (Vim-Ti and Vim-S) are shown in Table I.
TABLE I: Convolutional stem for patch embedding.
@@ IV-B Patch Embedding
The patch embedding in Vim can be implemented through a $16\times 16$ convolutional layer with a stride of 16. The convolutional stem for patch embedding in Vim-F comprises more convolutional layers, but the additional computational cost is negligible due to the use of smaller convolutional kernels. Considering Vim-Ti as the baseline model, the computational load measured in GFLOPs only increases by 0.035.
IV-C Vim-F Block
As mentioned earlier, the Vim-F block is an improvement over the Vim block, as shown in Fig. 3. Specifically, a 2D FFT is performed on the input to obtain the spectrogram. Then, the spectrogram is added to the original input, with the intensity of the frequency and spatial domain information being adjusted by two trainable parameters, $\alpha$ and $\beta$, respectively. After the linear layer mixes the hidden dimensions of the feature map, the cross-domain information is sent to the Vim block for processing.
Figure 3: Vim-F block
Vim-F requires a 2D discrete Fourier transform (DFT) to be performed on the feature map before scanning, thus introducing additional computation. The Cooley-Tukey FFT algorithm decomposes a DFT of a sequence with length N into two DFTs of sub-sequences with length N/2, significantly reducing the computational complexity. During the execution of the algorithm, the data is divided into two parts based on whether they are odd or even, and the DFTs of each part are computed separately before merging the results. For a 2D-DFT with dimensions $(L,D)$, its complexity is $\Omega(L\cdot D\cdot logL\cdot logD)$. The feedforward network is implemented by a linear layer, with a complexity of $\Omega(L\cdot D^{2})$. Therefore, compared to the Vim block, the additional overhead of the Vim-F block is approximately linearly related to the sequence length $L$.
IV-D Vim-F(H) Block
Observing Eq. (1), it is evident that since all elements of matrix $A$ are strictly confined within the range of 0 to 1 [5], the early information stored in the hidden state $h$ will continuously decay. This leads to Mamba being sensitive to the order of the input sequence. We propose that by weighting the input sequence with attention mechanisms before processing with Mamba, this strong local bias can be mitigated, thereby enhancing Mamba’s capability to handle long sequences. We integrate the Vim-F block with the linear attention [11] to maintain approximately linear complexity, which we term Vim-F(H), and its structure is depicted in Fig. 4.
Figure 4: Vim-F(H) block
TABLE II: Comparison of different backbones on ImageNet-1K classification.